OpenStudy (anonymous):

find the indefinite integral x^-13 dx

OpenStudy (anonymous):

when you are integrating a power series you add 1 to the exponent and divide by that exponent, and make to to add C since its indefinite

OpenStudy (anonymous):

\[\bf \int\limits_{}^{}x^ndx=\frac{ x^{n+1} }{ n+1 }\]\[\bf \int\limits_{}^{}x^{-13}dx=? \]

OpenStudy (anonymous):

+ C; I forgot the constant.

OpenStudy (anonymous):

x^-12/-12+C

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Awesome!! Thank you!!

OpenStudy (anonymous):

$$\begin{align*}\int x^{-13}\,\mathrm{d}x&=-\frac1{12}\int\left(-12x^{-13}\right)\,\mathrm{d}x\\&=-\frac1{12}\int\frac{\mathrm{d}}{\mathrm{d}x}\left(x^{-12}\right)\,\mathrm{d}x\\&=-\frac1{12}x^{-12}+C\end{align*}$$

OpenStudy (anonymous):

@oldrin.bataku Don't give away answers.

OpenStudy (anonymous):

It's not quite giving away an answer if it was already previously posted, not to mention I accompanied it with work that demonstrate why the rule works.$$\int x^n\,dx=\frac1{n+1}\int((n+1)x^n)\,dx=\frac1{n+1}\int\frac{d}{dx}\left(x^{n+1}\right)\,dx=\frac1{n+1}x^{n+1}+C$$

OpenStudy (anonymous):

The rule works because you are inversely differentiating..If:\[\bf f(x)=x^n \ then \ f'(x)=nx^{n-1}\], then inversely differentiating, i.e. adding 1 to the exponent and then dividing by this new exponent brings you to the antiderivative.

OpenStudy (anonymous):

This is obviously not an argument, just a complement to what you've already posted.